# A Survey of Neural Code Intelligence: Paradigms, Advances and Beyond

### v3

@@ Abstract
Neural Code Intelligence – leveraging deep learning to understand, generate, and optimize code – holds immense potential for transformative impacts on the whole society. Bridging the gap between Natural Language and Programming Language, this domain has drawn significant attention from researchers in both research communities over the past few years. This survey presents a systematic and chronological review of the advancements in code intelligence, encompassing over 50 representative models and their variants, more than 20 categories of tasks, and over 690 related works. We follow the historical progression to trace the paradigm shifts across different research phases (e.g., from modeling code with recurrent neural networks to the era of Large Language Models). Concurrently, we highlight the major technical transitions in models, tasks, and evaluations spanning through different stages. For applications, we also observe a co-evolving shift. It spans from initial endeavors to tackling specific scenarios, through exploring a diverse array of tasks during its rapid expansion, to currently focusing on tackling increasingly complex and varied real-world challenges. Building on our examination of the developmental trajectories, we further investigate the emerging synergies between code intelligence and broader machine intelligence, uncovering new cross-domain opportunities and illustrating the substantial influence of code intelligence across various domains. Finally, we delve into both the opportunities and challenges associated with this field, alongside elucidating our insights on the most promising research directions. An ongoing, dynamically updated project and resources associated with this survey have been released at https://github.com/QiushiSun/NCISurvey.
@@ 1 Introduction
Aware that the development of code intelligence is significantly reflected in the evolution of language models designed for code, we present in Figure 2 a chronological summary of representative works that outline the developmental trajectory of neural code intelligence. It provides a framework for this paper by outlining an overview of the technical advances within the domain. Building upon this timeline, we embrace a new perspective characterized by the paradigm shifts in models, applications, evaluations, and beyond, to explore these exciting advancements in code intelligence. We thoroughly investigate the literature and distill the key findings, techniques, and interconnections between research across various epochs. Moreover, we broaden our scope to explore the integration of other domains with code intelligence, discussing how code generation assists in machine reasoning, how code training enhances models’ mathematics capabilities, and how code serves as a medium to provide new approaches for solving typical NLP tasks. Furthermore, we explore a wide array of real-world cross-domain applications, spanning coding assistants, data science, autonomous agents, and AI4science. A GitHub project associated with this survey is actively maintained at https://github.com/QiushiSun/NCISurvey, which contains all the resources used to construct this paper, as well as comprehensive reading lists to assist readers in further exploration. We hope that this will facilitate the continued development of the field.
@@ 3.1.2 Encoder-Decoder
Meanwhile, SPT-Code [203] enhances its input by integrating linearized ASTs, thereby enabling the use of both natural language and code structures as inputs during the pre-training phase. To improve the code generation ability of T5-based models, researchers also explore strategies to enable the decoder part to learn syntax and data flow [204]. NatGen [195] represents an extension of CodeT5 that leverages the bimodal and dual-channel nature of structural information. Like DOBF, it is trained by “Naturalizing” source code to exploit the codes’ naturalness and semantics. It requires the model to receive “unnatural” synthetic code as input and produce semantically equivalent code, mirroring the quality and style a human developer would prefer to write. CodeT5Mix [205] is composed of a mixture of encodes and decoders, each with specific code functionality. They can be flexibly combined to suit different scenarios and enjoy mutual benefits from joint pretraining on various targets. Further, weight-sharing strategies in decoders are used to act as task-specific experts to reduce interference across code-related tasks. Very recently, AST-T5 [206] employs a structure-aware code segmentation method during its training process, enabling the model to reconstruct code structures at various granularities.
@@ 4.1 Large Language Models for Code
In terms of application paradigms, there is no longer a necessity to meticulously select annotated data for training from scratch or to engage in task-specific fine-tuning. As illustrated in Figure 4, the primary approach now leans towards leveraging prompt learning and providing relevant in-context demonstrations, which are usually non-invasive to models [4, 279]. Consequently, the approach to handling various code-related tasks has gradually evolved from the diversified forms mentioned in Section 2.2 to predominantly generative methods [280].
@@
Regarding CodeLLMs that are built from additional pre-training, CodeGeeX2 [290] represents the second iteration of the CodeGeeX model lineage. Unlike its predecessor, which was trained from scratch, Diverging from its predecessor trained from scratch, it is developed upon ChatGLM2 [331, 332] with further pre-training on code tokens. Similarly, Code-Qwen [297] follows a training approach akin to CodeLLaMA [294], using base model Qwen trained on a combination of text and code data as an initialization and then continuing to pre-train on code data. This approach is mirrored in other models such as AquilaCode [^12] https://huggingface.co/BAAI/AquilaCode-multi, which also enhances base models with extra training on code corpora. Recently released CodeGemma [305] denotes a compilation of lightweight models crafted through further code infilling training on Gemma [333], particularly adept at code completion and generation from provided code prefixes or suffixes.
@@ 4.2 Learning with Execution Feedback
Another pathway to further enhance CodeLLMs involves integrating Reinforcement Learning (RL), which incorporates non-differentiable reward signals into the training process. Unlike approaches represented by reinforcement learning from human feedback (RLHF) that utilize human preferences [343, 282, 344], the inherently compilable and executable nature of codes allows for the use of compilers or interpreters to automatically generate precise feedback. Such endeavors were initiated in the era of CodePTMs, as COMPCODER [345] harnesses the compilability signals to optimize both the generator (e.g., CodeGPT [114]) and the discriminator (MLPs) via RL strategies.
@@ 4.3.1 The Shift in Evaluation Metrics
This revised method involves producing $n\geq k$ samples for each task, counting correct samples $c\leq n$ that pass unit tests successfully, thereby deriving an unbiased estimator. pass@$k$ represents a milestone in NL2Code evaluation, which transcends the limitations of earlier metrics incapable of accurately reflecting functional correctness and has become a standard practice in all subsequent benchmarks.
@@ 6.1 Boosting Software Develop Workflows
Serving as Coding Assistants. It is an indisputable fact that coding assistants, exemplified by GitHub Copilot [453], are fundamentally transforming the landscape of software development [454]. Earlier attempts, such as aiXcoder [^18] https://github.com/aixcoder-plugin and Intellicode [198], have already demonstrated that deep learning models can assist with basic coding tasks. With the advent of the LLM era, emerging products like Code Whisperer [455], Tabnine [456], and Coze [^19] https://coze.com/, which are built on commercial models, have begun to mature into established products. Concurrently, coding assistants based on open-source CodeLLMs are increasingly gaining popularity and are being utilized in multiple code downstream tasks. For instance, FauxPilot [^20] https://github.com/fauxpilot/fauxpilot and Starchat [308], respectively based on the CodeGen and BigCode models, can serve as locally hosted alternatives to Copilot. These versatile tools can be used for code generation, completion, repair, and even predicting the time complexity [457]. Additionally, when paired with a code interpreter (e.g., ChatGPT Plugins [^21] https://openai.com/blog/chatgpt-plugins#code-interpreter), chatbots are capable of aiding users in building temporary programs within conversations. Furthermore, interacting with users in the form of extensions has also started to become popular, with CodeGeeX [290] already adapting itself to IDEs such as Visual Studio Code and JetBrains. Empirical research has illuminated the profound impact of these AI coding tools on developer efficiency. For instance, A prior study [458] has indicated that the majority of GitHub Copilot users have experienced enhanced programming efficiency and improved code quality when coding with it. Complementarily, Peng et al. [459] quantify this enhancement in productivity, demonstrating a notable acceleration in task completion when developers employ Copilot.
@@
As for leveraging code intelligence within GitHub, CodeAgent [476] suggests the use of external tools (e.g., Format Checker) for repo-level code generation. Meanwhile, beyond using pre-defined function set [477, 478], GitAgent [479] autonomously integrates repositories in response to user queries, thereby expanding its toolkit. Regarding software maintenance, the quality of code documentation is directly linked to development efficiency [480]. To automate this, RepoAgent [481] has been proposed for documentation generation. It utilizes AST analysis to understand code structure and discern reference relationships within files, providing a contextual perspective for LLMs to assist in identifying the functional semantics to support fine-grained code documentation generation. For more reliable testing, Meta has introduced TestGen-LLM [482], which aims to improve existing human-written tests for automated unit test generation. The majority of its improvements have landed in industrial production. Meanwhile, Google has advocated for the automation of resolving review comments within daily development workflows [483].
@@ 6.3 Building Code-Empowered Agents
A perennial topic in machine intelligence is to develop agent systems [504], such as robots, that can perform complex tasks requiring interaction with real-world environments. Recent breakthroughs in LLMs have notably enhanced the task-solving capabilities of AI agents. Typically focused on natural language generation, LLMs offer immense potential [326, 505, 325], yet the inherent ambiguity of natural language can sometimes render the precision and efficiency of planning and interaction [506]. To address this, emerging research advocates the integration of code as the de facto standard of agent systems [507, 508, 509, 510, 511]. The following discussions will delve into these advancements.
@@ 7.3 Repo-Level Code Understanding and Generation
Following the surge in the popularity of CodeLLMs, their applications have primarily focused on individual functions or files. However, in real-world software engineering, developers often need to consider the relationships between different files and functions within a code repository. When extending the tasks discussed in Section 2.2 to broader scenarios, such as completing or repairing a code snippet within a repository [581], The capability of models to invoke variables and functions from other files remains underexplored. This may necessitate the integration of code understanding, cross-file dependencies understanding, and retrieval-augmented generation [582]. Also, given the potential complexity of code repositories, this may require research into how to expand CodeLLMs’ efficient context window [583] and their ability to understand cross-file dependencies. Therefore, repository-level code generation and understanding represent a direction of both practical and research significance, with production environments also serving as a sustainable and challenging testbed for future models.
@@ A.1 Reading Lists
In the NCISurvey project related to this paper, we have prepared a curated reading list for our readers, encompassing a diverse array of research areas within the realm of code intelligence.
@@ E.2 Publication Trends Data and Milestones
The construction of Figure 1 was achieved through the use of the ArXiv advanced search [^23] https://arxiv.org/search/advanced, with the calculation based on an exact match by querying a group of keywords (e.g., code representation, code generation) in the title or abstract. Detailed information on this is provided within the NCISurvey project.
@@ Acknowledgement
Purpose Dataset Date # PLs. Description Eval. Metric Open Domain CONCODE [159] [link] 2018 1 a dataset with 100K examples consisting of Java classes from online code repositories EM/BLEU ODEX [353] [link] 2023 1 an open-domain execution-based natural language to Python code generation dataset pass@k Code Exercise HumanEval [36] [link] 2021 1 a dataset of 164 handwritten programming problems with unit tests pass@k MBPP [354] [link] 2021 1 a dataset containing 974 short Python programs pass@k MathQA-Python [354] [link] 2021 1 a Python version of the MathQA benchmark contains 23K problems Acc. AixBench [682] [link] 2022 1 a code generation benchmark dataset Pass Rate BIG-Bench [355, 328] [link] 2023 - a benchmark containing over 12 tasks can be solved by coding - CoderEval [683] [link] 2023 2 a dataset consisting of 230 Python and 230 Java code generation tasks Pass Rate CodeApex [584] [link] 2023 1 a bilingual programming evaluation benchmark for large language models Acc. ClassEval [684] [link] 2023 1 A handcrafted benchmark for evaluating class-level code generation pass@k OOP [685] [link] 2024 1 an OOP-focused benchmark, with 431 Python programs that encompass OOP concepts like classes pass@k NCB [686] [link] 2024 2 402 high-quality python and java problems, selected from natural user queries from coding services pass@k Competitions APPS [356] [link] 2021 1 a benchmark including 10K problems for Python code generation Acc. CodeContests [273] [link] 2022 3 a dataset of competitive programming problems pass@k TACO [687] [link] 2023 1 a dataset on algorithmic code generation, consisting of 25K training problems and 1K testing problem pass@k Multilingual MultiPL-E [161] [link] 2022 18 a parallel, multilanguage benchmark for natural-language-to-code generation pass@k MCoNaLa [688] [link] 2022 1 a benchmark for code generation from multiple natural languages BLEU MBXP [357] [link] 2023 12 a benchmark to evaluate code generation models in over 10 programming languages pass@k MathQA-X [357] [link] 2023 3 a dataset transpiled from the original Python dataset MathQA Pass Rate/Acc. xCodeEval [654] [link] 2023 11 an executable dataset of 450K small buggy/fixed program pairs Pass Rate HumanEval-X [290] [link] 2023 4 a benchmark of 164 code problems for evaluating multilingual models pass@k Data Science JuICe [162] [link] 2018 1 a corpus of 1.5M examples with a curated test set of 3.7K instances EM/BLEU PlotCoder [689] [link] 2021 1 a dataset of plot samples extracted from the original dev and test splits of JuICe Acc. DSP [163] [link] 2022 1 a collection of 1.1K problems curated from 306 pedagogical notebooks pass@k ExeDS [690] [link] 2022 1 a dataset of 534 problems for execution evaluation for data science code generation tasks (Code)BLEU/EM DS-1000 [329] [link] 2023 1 a code generation benchmark with 1K data science problems spanning 7 Python libraries pass@k DAEval [691] [link] 2024 1 a dataset comprising 257 data analysis questions derived from 52 CSV files Pass Rate Python Libs PandasEval [199] [link] 2022 1 a dataset consisting of 101 programming problems on Pandas library pass@k NumpyEval [199] [link] 2022 1 a dataset consisting of 101 programming problems on Numpy library pass@k TorchDataEval [358] [link] 2022 1 a dataset with 50 programming problems using the TorchData library pass@k Multi-Turn MTPB [275] [link] 2023 1 an open benchmark consisting of 115 diverse problem sets that are factorized into multi-turn prompts Pass Rate Command Line NL2Bash [160] [link] 2018 1 a corpus of 9K English-command pairs, covering over 100 unique Bash utilities Acc. AI4Science BioCoder [359] [link] 2023 2 a benchmark developed to evaluate large language models in generating bioinformatics-specific code pass@k Education StudentEval [692] [link] 2023 1 a benchmark containing 1.7K prompts for 48 problems, written by 80 students pass@1 COJ2022 [87] [link] 2023 1 a dataset containing 5.9K C programs with semantic errors submitted to 498 different assignments F1 Log Parsing LogHub1 [693] [link] 2019 - a dataset of logs produced by 16 different systems Acc. LogHub2[694] [link] 2023 - a suite of 14 datasets, each averaging 3.6M log lines FGA Hardware Design VerilogEval [695] [link] 2023 1 a dataset consists of 156 problems from the Verilog instructional website HDLBits BLEU/pass@k

### v4

@@ Abstract
Neural Code Intelligence – leveraging deep learning to understand, generate, and optimize code – holds immense potential for transformative impacts on the whole society. Bridging the gap between Natural Language and Programming Language, this domain has drawn significant attention from researchers in both research communities over the past few years. This survey presents a systematic and chronological review of the advancements in code intelligence, encompassing over 50 representative models and their variants, more than 20 categories of tasks, and over 700 related works. We follow the historical progression to trace the paradigm shifts across different research phases (e.g., from modeling code with recurrent neural networks to the era of Large Language Models). Concurrently, we highlight the major technical transitions in models, tasks, and evaluations spanning through different stages. For applications, we also observe a co-evolving shift. It spans from initial endeavors to tackling specific scenarios, through exploring a diverse array of tasks during its rapid expansion, to currently focusing on tackling increasingly complex and varied real-world challenges. Building on our examination of the developmental trajectories, we further investigate the emerging synergies between code intelligence and broader machine intelligence, uncovering new cross-domain opportunities and illustrating the substantial influence of code intelligence across various domains. Finally, we delve into both the opportunities and challenges associated with this field, alongside elucidating our insights on the most promising research directions. An ongoing, dynamically updated project and resources associated with this survey have been released at https://github.com/QiushiSun/Awesome-Code-Intelligence.
@@ 1 Introduction
Aware that the development of code intelligence is significantly reflected in the evolution of language models designed for code, we present in Figure 2 a chronological summary of representative works that outline the developmental trajectory of neural code intelligence. It provides a framework for this paper by outlining an overview of the technical advances within the domain. Building upon this timeline, we embrace a new perspective characterized by the paradigm shifts in models, applications, evaluations, and beyond, to explore these exciting advancements in code intelligence. We thoroughly investigate the literature and distill the key findings, techniques, and interconnections between research across various epochs. Moreover, we broaden our scope to explore the integration of other domains with code intelligence, discussing how code generation assists in machine reasoning, how code training enhances models’ mathematics capabilities, and how code serves as a medium to provide new approaches for solving typical NLP tasks. Furthermore, we explore a wide array of real-world cross-domain applications, spanning coding assistants, data science, autonomous agents, and AI4science. A GitHub project associated with this survey is actively maintained at https://github.com/QiushiSun/Awesome-Code-Intelligence, which contains the resources used to construct this paper, as well as comprehensive reading lists to support further exploration. We hope that this will facilitate the continued development of the field.
@@ 3.1.2 Encoder-Decoder
Meanwhile, SPT-Code [203] enhances its input by integrating linearized ASTs, thereby enabling the use of both natural language and code structures as inputs during the pre-training phase. To improve the code generation ability of T5-based models, researchers also explore strategies to enable the decoder part to learn syntax and data flow [204]. NatGen [195] represents an extension of CodeT5 that leverages the bimodal and dual-channel nature of structural information. Like DOBF, it is trained by “Naturalizing” source code to exploit the codes’ naturalness and semantics. It requires the model to receive “unnatural” synthetic code as input and produce semantically equivalent code, mirroring the quality and style a human developer would prefer to write. CodeT5Mix [205] is composed of a mixture of encoders and decoders, each with specific code functionality. They can be flexibly combined to suit different scenarios and enjoy mutual benefits from joint pretraining on various targets. Further, weight-sharing strategies in decoders are used to act as task-specific experts to reduce interference across code-related tasks. Very recently, AST-T5 [206] employs a structure-aware code segmentation method during its training process, enabling the model to reconstruct code structures at various granularities.
@@ 4.1 Large Language Models for Code
In terms of application paradigms, there is no longer a necessity to meticulously select annotated data for training from scratch or to engage in task-specific fine-tuning. As illustrated in Figure 4, the primary approach now leans towards leveraging prompt learning and providing relevant in-context demonstrations [279], which are usually non-invasive to models [4, 280]. Consequently, the approach to handling various code-related tasks has gradually evolved from the diversified forms mentioned in Section 2.2 to predominantly generative methods [281].
@@
Regarding CodeLLMs that are built from additional pre-training, CodeGeeX2 [291] represents the second iteration of the CodeGeeX model lineage. [(zhirui): Semantic repetition] Unlike its predecessor, which was trained from scratch, Diverging from its predecessor trained from scratch, it is developed upon ChatGLM2 [332, 333] with further pre-training on code tokens. Similarly, Code-Qwen [298] follows a training approach akin to CodeLLaMA [295], using base model Qwen trained on a combination of text and code data as an initialization and then continuing to pre-train on code data. This approach is mirrored in other models such as AquilaCode [^12] https://huggingface.co/BAAI/AquilaCode-multi, which also enhances base models with extra training on code corpora. Recently released CodeGemma [306] denotes a compilation of lightweight models crafted through further code infilling training on Gemma [334], particularly adept at code completion and generation from provided code prefixes or suffixes.
@@ 4.2 Learning with Execution Feedback
Another pathway to further enhance CodeLLMs involves integrating Reinforcement Learning (RL), which incorporates non-differentiable reward signals into the training process. Unlike approaches represented by reinforcement learning from human feedback (RLHF) that utilize human preferences [344, 283, 345], The inherently compilable and executable nature of codes allows compilers or interpreters to automatically generate precise feedback. Such endeavors were initiated in the era of CodePTMs, as COMPCODER [346] harnesses the compilability signals to optimize both the generator (e.g., CodeGPT [114]) and the discriminator (MLPs) via RL strategies.
@@ 4.3.1 The Shift in Evaluation Metrics
This revised method involves producing $n\geq k$ samples for each task, counting correct samples $c\leq n$ that pass unit tests successfully, thereby deriving an unbiased estimator. pass@$k$ represents a milestone in NL2Code evaluation, which transcends the limitations of earlier metrics incapable of accurately reflecting functional correctness and has become a standard practice in all subsequent benchmarks. Building on this, contemporary work has also started incorporating elements like self-debugging and candidate generation into the evaluation pipeline [369].
@@ 6.1 Boosting Software Develop Workflows
Serving as Coding Assistants. It is an indisputable fact that coding assistants, exemplified by GitHub Copilot [455], are fundamentally transforming the landscape of software development [456]. Earlier attempts, such as aiXcoder [^18] https://github.com/aixcoder-plugin and Intellicode [198], have already demonstrated that deep learning models can assist with basic coding tasks. With the advent of the LLM era, emerging products like Code Whisperer [457], Tabnine [458], Coze [^19] https://coze.com/, and aiXcoder [459].
which are built on commercial models, have begun to mature into established products. Concurrently, coding assistants based on open-source CodeLLMs are increasingly gaining popularity and are being utilized in multiple code downstream tasks. For instance, FauxPilot [^20] https://github.com/fauxpilot/fauxpilot and Starchat [309], respectively based on the CodeGen and BigCode models, can serve as locally hosted alternatives to Copilot. These versatile tools can be used for code generation, completion, repair, and even predicting the time complexity [460]. Additionally, when paired with a code interpreter (e.g., ChatGPT Plugins [^21] https://openai.com/blog/chatgpt-plugins#code-interpreter), chatbots are capable of aiding users in building temporary programs within conversations. Furthermore, interacting with users in the form of extensions has also started to become popular, with CodeGeeX [291] already adapting itself to IDEs such as Visual Studio Code and JetBrains. Empirical research has illuminated the profound impact of these AI coding tools on developer efficiency. For instance, A prior study [461] has indicated that the majority of GitHub Copilot users have experienced enhanced programming efficiency and improved code quality when coding with it. Complementarily, Peng et al. [462] quantify this enhancement in productivity, demonstrating a notable acceleration in task completion when developers employ Copilot.
@@
As for leveraging code intelligence within GitHub, CodeAgent [479] suggests the use of external tools (e.g., Format Checker) for repo-level code generation. Meanwhile, beyond using pre-defined function set [480, 481], GitAgent [482] autonomously integrates repositories in response to user queries, thereby expanding its toolkit. Regarding software maintenance, the quality of code documentation is directly linked to development efficiency [483]. To automate this, RepoAgent [484] has been proposed for documentation generation. It utilizes AST analysis to understand code structure and discern reference relationships within files, providing a contextual perspective for LLMs to assist in identifying the functional semantics to support fine-grained code documentation generation. For more reliable testing, Meta has introduced TestGen-LLM [485], which aims to improve existing human-written tests for automated unit test generation. The majority of its improvements have landed in industrial production. Meanwhile, Google has advocated for the automation of resolving review comments within daily development workflows [486]. Recently released MarsCode Agent [^22] https://www.marscode.com/ integrates AI-powered tools into a cloud-based development environment, streamlining coding processes with features like code completion, explanation, and automated bug fixing.
@@ 6.3 Building Code-Empowered Agents
A perennial topic in machine intelligence is to develop agent systems [507], such as robots, that can perform complex tasks requiring interaction with real-world environments. Recent breakthroughs in LLMs have notably enhanced the task-solving capabilities of AI agents [508]. Typically focused on natural language generation, LLMs offer immense potential [327, 509, 326], yet the inherent ambiguity of natural language can sometimes render the precision and efficiency of planning and interaction [510]. To address this, emerging research advocates the integration of code as the de facto standard of agent systems [511, 512, 513, 514, 515]. The following discussions will delve into these advancements.
@@ 7.3 Repo-Level Code Understanding and Generation
Following the surge in the popularity of CodeLLMs, their applications have primarily focused on individual functions or files. However, in real-world software engineering, developers often need to consider the relationships between different files and functions within a code repository. When extending the tasks discussed in Section 2.2 to broader scenarios, such as completing or repairing a code snippet within a repository [585, 479], The capability of models to invoke variables and functions from other files remains underexplored. This may necessitate the integration of code understanding, cross-file dependencies understanding, and retrieval-augmented generation [586]. Also, given the potential complexity of code repositories, this may require research into how to expand CodeLLMs’ efficient context window [587] and their ability to understand cross-file dependencies. Therefore, repository-level code generation and understanding represent a direction of both practical and research significance, with production environments also serving as a sustainable and challenging testbed for future models.
@@ A.1 Reading Lists
In the Awesome-Code-Intelligence project related to this paper, we have prepared a curated reading list for our readers, encompassing a diverse array of research areas within the realm of code intelligence.
@@ E.2 Publication Trends Data and Milestones
The construction of Figure 1 was achieved through the use of the ArXiv advanced search [^24] https://arxiv.org/search/advanced, with the calculation based on an exact match by querying a group of keywords (e.g., code representation, code generation) in the title or abstract. Detailed information on this is provided within the Awesome-Code-Intelligence project.
@@ Acknowledgement
Author Contributions
Writing this comprehensive survey and continuously updating its contents is no easy job. We are deeply grateful to the authors for their support and dedication to this work. Their contributions to this paper are as follows:
  • Project concept and leadership: Qiushi Sun, Xiang Li, and Zhiyong Wu.
  • Paper Writing: Qiushi Sun, Fangzhi Xu, Chang Ma, Kanzhi Cheng, Zhirui Chen, Qipeng Guo, and Lingpeng Kong.
  • Paper Revising: Qiushi Sun, Chang Ma, Zhirui Chen, Chengcheng Han, Renyu Zhu, Lingpeng Kong, Fei Yuan, Qipeng Guo, Pengcheng Yin, Xipeng Qiu, Xiang Li, and Xiaoli Li.
  • Experiments: Fangzhi Xu and Qiushi Sun.
  • GitHub project maintenance: Shuai Yuan, Zhirui Chen, Qiushi Sun, Jianing Wang, Kanzhi Cheng, Fangzhi Xu, and Zhangyue Yin.
  • Strategic advice: Pengcheng Yin, Xipeng Qiu Xiang Li, and Xiaoli Li.
@@
  • Update on October 25, 2024: add more related models, and updates on coding agents.
  • Update on November 1, 2024: add more related models, benchmarks, methods, and author contribution list.
@@
Purpose Dataset Date # PLs. Description Eval. Metric Open Domain CONCODE [159] [link] 2018 1 a dataset with 100K examples consisting of Java classes from online code repositories EM/BLEU ODEX [354] [link] 2023 1 an open-domain execution-based natural language to Python code generation dataset pass@k Code Exercise HumanEval [36] [link] 2021 1 a dataset of 164 handwritten programming problems with unit tests pass@k MBPP [355] [link] 2021 1 a dataset containing 974 short Python programs pass@k MathQA-Python [355] [link] 2021 1 a Python version of the MathQA benchmark contains 23K problems Acc. AixBench [686] [link] 2022 1 a code generation benchmark dataset Pass Rate BIG-Bench [356, 329] [link] 2023 - a benchmark containing over 12 tasks can be solved by coding - CoderEval [687] [link] 2023 2 a dataset consisting of 230 Python and 230 Java code generation tasks Pass Rate CodeApex [588] [link] 2023 1 a bilingual programming evaluation benchmark for large language models Acc. ClassEval [688] [link] 2023 1 a handcrafted benchmark for evaluating class-level code generation pass@k OOP [689] [link] 2024 1 an OOP-focused benchmark, with 431 Python programs that encompass OOP concepts like classes pass@k NCB [690] [link] 2024 2 402 high-quality python and java problems, selected from natural user queries from coding services pass@k MHPP [691] [link] 2024 1 a dataset of 140 Python problems designed to assess LLMs’ advanced code generation abilities. pass@k Competitions APPS [357] [link] 2021 1 a benchmark including 10K problems for Python code generation Acc. CodeContests [273] [link] 2022 3 a dataset of competitive programming problems pass@k TACO [692] [link] 2023 1 a dataset on algorithmic code generation, consisting of 25K training problems and 1K testing problem pass@k Multilingual MultiPL-E [161] [link] 2022 18 a parallel, multilanguage benchmark for natural-language-to-code generation pass@k MCoNaLa [693] [link] 2022 1 a benchmark for code generation from multiple natural languages BLEU MBXP [358] [link] 2023 12 a benchmark to evaluate code generation models in over 10 programming languages pass@k MathQA-X [358] [link] 2023 3 a dataset transpiled from the original Python dataset MathQA Pass Rate/Acc. xCodeEval [658] [link] 2023 11 an executable dataset of 450K small buggy/fixed program pairs Pass Rate HumanEval-X [291] [link] 2023 4 a benchmark of 164 code problems for evaluating multilingual models pass@k Data Science JuICe [162] [link] 2018 1 a corpus of 1.5M examples with a curated test set of 3.7K instances EM/BLEU PlotCoder [694] [link] 2021 1 a dataset of plot samples extracted from the original dev and test splits of JuICe Acc. DSP [163] [link] 2022 1 a collection of 1.1K problems curated from 306 pedagogical notebooks pass@k ExeDS [695] [link] 2022 1 a dataset of 534 problems for execution evaluation for data science code generation tasks (Code)BLEU/EM DS-1000 [330] [link] 2023 1 a code generation benchmark with 1K data science problems spanning 7 Python libraries pass@k DAEval [696] [link] 2024 1 a dataset comprising 257 data analysis questions derived from 52 CSV files Pass Rate Python Libs PandasEval [199] [link] 2022 1 a dataset consisting of 101 programming problems on Pandas library pass@k NumpyEval [199] [link] 2022 1 a dataset consisting of 101 programming problems on Numpy library pass@k TorchDataEval [359] [link] 2022 1 a dataset with 50 programming problems using the TorchData library pass@k Multi-Turn MTPB [275] [link] 2023 1 an open benchmark consisting of 115 diverse problem sets that are factorized into multi-turn prompts Pass Rate Command Line NL2Bash [160] [link] 2018 1 a corpus of 9K English-command pairs, covering over 100 unique Bash utilities Acc. AI4Science BioCoder [360] [link] 2023 2 a benchmark developed to evaluate large language models in generating bioinformatics-specific code pass@k SciCode [697] 2024 1 a scientist-curated benchmark containing 80 real scientific coding problems pass@k Education StudentEval [698] [link] 2023 1 a benchmark containing 1.7K prompts for 48 problems, written by 80 students pass@1 COJ2022 [87] [link] 2023 1 a dataset containing 5.9K C programs with semantic errors submitted to 498 different assignments F1 Log Parsing LogHub1 [699] [link] 2019 - a dataset of logs produced by 16 different systems Acc. LogHub2[700] [link] 2023 - a suite of 14 datasets, each averaging 3.6M log lines FGA Hardware Design VerilogEval [701] [link] 2023 1 a dataset consists of 156 problems from the Verilog instructional website HDLBits BLEU/pass@k